# Multi-Path Collaborative Reasoning via Reinforcement Learning

### v1

@@ 1 Introduction
The advent of Chain-of-Thought (CoT) [51, 21, 14] has markedly improved the reasoning capabilities of Large Language Models (LLMs) [29, 50, 43] on complex tasks. By generating explicit intermediate steps in natural language, CoT decomposes problems and guides models step-by-step toward the answer [4]. However, the discrete token decoding underlying standard CoT induces internal determinism, which limits the exploration of plausible alternatives [57].
Recent efforts seek to transcend discrete token decoding by introducing continuity into the reasoning process. Latent reasoning [15, 39, 41] propagates hidden states instead of tokens to enable differentiable reasoning, but such representations often deviate from the pretrained embedding space, requiring costly alignment and limiting compatibility [38]. To address this, Soft Thinking [61, 52] performs soft aggregation within the input embedding space, preserving model compatibility while simulating a smooth semantic flow.
Nevertheless, we reveal that Soft Thinking inherently lacks the ability to represent diverse semantic trajectories in parallel [52]. As shown in Figure  2, during consecutive decoding, each subsequent distribution consistently follows the dominant semantic token from the previous step, resulting in a coherent but exclusive progression along a single trajectory. Although soft aggregation improves information capacity, it primarily reinforces the dominant path while introducing extraneous noise over time, as evidenced in Figures 1(c) and 1(d).
This phenomenon stems from the greedy nature of LLMs in autoregressive generation, where hidden states evolve at each step toward the most confident semantic direction [52]. Such locally optimal decisions are amplified over long reasoning chains, resulting in a cumulative bias toward dominant paths and exclusion of alternatives. Thus, when a reasoning trajectory begins with a flawed premise, the lack of timely corrective feedback allows erroneous logic to propagate unimpeded through subsequent steps [53].
@@
This paradigm promotes robust policy optimization by allowing trajectories to benefit from cross-path insights. Specifically, the parameter-free gate facilitates step-level information exchange, providing trajectories with opportunities to gently refine their reasoning chain and reduce the persistence of local biases. In particular, by maintaining and refining multiple hypotheses in parallel, M3PO aligns with human-like cognition [10, 9, 2], striking an adaptive balance between exploration and exploitation.
@@ 2.1 Continuous Space Reasoning
Recent works have explored continuous space reasoning as an alternative to explicit CoT generation [31, 12, 49, 46, 45]. Several studies demonstrate that Transformers can perform multi-hop reasoning implicitly through their internal hidden states, without generating intermediate textual steps [55, 36]. Deng et al. [7, 8] further exploit this idea by using hidden dynamics directly for reasoning, bypassing symbolic output altogether. Geiping et al. [11] extend this paradigm with a depth-recurrent architecture that reuses Transformer layers during inference, increasing computational depth per token and enabling internal processing.
Another line focuses on explicit continuous reasoning through latent representations [7, 41]. COCONUT [15] enables reasoning in the hidden state space by replacing discrete tokens with continuous vectors, eliminating the need for human-readable chains. CODI [39] learns to align recurrent state updates via self-distillation, forming structured reasoning trajectories in the latent space. Although conceptual advances, they often struggle to scale to larger models and maintain performance on complex reasoning tasks.
@@
More recently, training-free approaches such as Soft Thinking [61, 13, 52] and Mixture-of-Inputs (MoI) [62] have emerged, leveraging output distributions to align hidden states with input embeddings, facilitating iterative reasoning in continuous space. Building on this, HRPO [59] introduces a hybrid policy that fuses discrete token embeddings with soft embeddings through a gating mechanism, enabling richer internal state evolution during reasoning.
@@ 2.2 Reinforcement Learning
Reinforcement learning (RL) has emerged as a powerful framework for enhancing language models through feedback-driven decision making. In natural language generation, RL enables models to learn from implicit human preferences by maximizing cumulative rewards [30]. This is typically achieved with policy gradient methods such as REINFORCE [40], which update the policy using sampled responses and their rewards. Actor-critic approaches like A2C [28] reduce variance with learned baselines, while PPO [34] improves stability through a clipped objective that limits policy changes and prevents performance collapse.
An alternative line of work uses preference-based optimization, where models are fine-tuned on pairwise response comparisons [35, 3, 60, 48]. Methods like DPO [33] bypass explicit reward modeling by directly optimizing for preferred outputs, often eliminating the need for a reference model. Yet despite their efficiency, offline approaches generally lag behind online RL in complex reasoning settings. Recent online methods such as RLOO [1], GRPO [37], and REINFORCE++ [20] address variance and scalability by computing baselines from multiple responses in the same batch. Leveraging group-level comparisons, they achieve more stable and memory-efficient training, making them well suited for long-horizon reasoning tasks.
@@ 3.3 Policy Optimization
Note that M3PO uses raw log probabilities in its policy gradient (Equation (8)) rather than likelihood ratios as in PPO [34] or GRPO [37], eliminating the need for ratio clipping under our conservative update schedule. Furthermore, each trajectory is used only once for gradient updates since the hybrid representations are intrinsically tied to $\theta$, maintaining strict on-policy training. This makes M3PO both lightweight and compatible with other RL optimizations.
@@ 4.1 Evaluation on Knowledge Benchmarks
For knowledge-intensive reasoning, we use five widely adopted open-domain and multi-hop QA datasets: Natural Questions (NQ) [24], TriviaQA [23], HotpotQA [56], 2WikiMultiHopQA (2WikiMQA) [18], and Bamboogle [32]. Following HRPO [59], we employ the E5-base embedding model [47] to retrieve the top three Wikipedia documents as context for each prompt. We train on a joint corpus formed by merging the NQ and HotpotQA training sets, and we report exact match results on each dataset’s official evaluation split. Detailed implementation settings are provided in the Appendix.
For the 1.5B and 3B model sizes [54], we evaluate M3PO against representative post-training and retrieval baselines, including Supervised Fine-Tuning (SFT), the RL methods PPO [34] and GRPO [37], Retrieval-Augmented Generation (RAG) [25], and Hybrid Reasoning Policy Optimization (HRPO) [59], an RL-based approach built on soft thinking [61]. We also report results with the larger Qwen2.5-7B-Instruct [54] backbone using standard QA and retrieval pipelines, namely direct inference (QA), Chain-of-Thought (CoT) [51], interleaving retrieval with CoT (IRCoT) [44], Search-o1 [26], and RAG [25]. The best result in each block of Table 1 is highlighted in bold.
@@ 4.2 Evaluation on STEM Benchmarks
For reasoning-intensive STEM evaluation, we employ five benchmarks spanning mathematical and scientific domains: GSM8k [6], MATH [17], MATH500 [27], MMLU-STEM [16], and ARC-Challenge [5]. For training, we use the GSM8k training split for GSM8k, the MATH training split for MATH and MATH500, and a merged corpus that combines the auxiliary MMLU and ARC-C training sets for MMLU-ST and ARC-C. Detailed settings are provided in Appendix.
For models at the 1.5B and 3B scales [54], we compare against several strong baselines: SFT, standard RL methods like PPO [34] and GRPO [37], and the RL-based latent reasoning method HRPO [59]. To further contextualize the performance of M3PO, we also include comparisons with larger models ($\geq$ 7B parameters) using few-shot CoT reasoning, specifically DeepSeekMath-7B [37], Gemma-2-9B [42], Qwen2.5-7B [54], and MAmmoTH2-7B [58].
@@ 6 Implementation
M3PO is a lightweight multi-path collaborative learning framework designed to be compatible with any LLM architecture. It introduces no additional trainable parameters and preserves the original inference efficiency of the base model. To enable efficient training under this framework, we integrate optimized kernel implementations from Unsloth [^1] https://github.com/unslothai/unsloth and apply low-rank adaptation (LoRA) [19] for parameter-efficient fine-tuning. All M3PO experiments adopt the hyperparameter settings detailed in Table 3, which serve as our default unless otherwise specified.
@@
The prompt construction follows a consistent template during both training and evaluation. Each input begins with a system message instructing the model to carry out step-by-step reasoning before generating its final answer, followed by the user query. The full prompt is then formatted according to the model’s native chat template. In line with established practice [59], we use the minimal delimiter #### to separate the model’s reasoning chain from its final answer. This delimiter tokenizes as a single unit, adding no sequence length overhead while providing a clear signal for transitioning from latent collaborative reasoning to autoregressive answer generation. To ensure reasoning integrity, a penalty mechanism assigns zero reward to completions containing repeated delimiters, regardless of answer correctness. This prevents early termination of the reasoning process and encourages complete reasoning chains. Full prompt examples for different task types, including system messages and representative queries, are provided in Figures 9, 10, and 11, respectively.
@@
We use greedy decoding and the standard inference pipeline of the base LLM to ensure reproducibility. For knowledge tasks, exact match scores are reported on validation and test splits following [22]. For mathematical reasoning benchmarks including GSM8K, MATH, and MATH-500, along with multiple-choice datasets (MMLU-STEM and ARC-Challenge), we follow the post-processing and scoring procedures defined in [58].

### v2

@@ 1 Introduction
The advent of Chain-of-Thought (CoT) [wei2022chain, jaech2024openai, guo2025deepseek] has markedly improved the reasoning capabilities of Large Language Models (LLMs) [openai2023gpt, wei2022emergent, touvron2023llama] on complex tasks. By generating explicit intermediate steps in natural language, CoT decomposes problems and guides models step-by-step toward the answer [cho2022advances]. However, the discrete token decoding underlying standard CoT induces internal determinism, which limits the exploration of plausible alternatives [yao2023tree].
Recent efforts seek to transcend discrete token decoding by introducing continuity into the reasoning process. Latent reasoning [hao2024training, shen2025codi, tack2025llm] propagates hidden states instead of tokens to enable differentiable reasoning, but such representations often deviate from the pretrained embedding space, requiring costly alignment and limiting compatibility [shen2025efficient]. To address this, Soft Thinking [zhang2025soft, wu2025llms] performs soft aggregation within the input embedding space, preserving model compatibility while simulating a smooth semantic flow.
Nevertheless, we reveal that Soft Thinking inherently lacks the ability to represent diverse semantic trajectories in parallel [wu2025llms]. As shown in Figure  2, during consecutive decoding, each subsequent distribution consistently follows the dominant semantic token from the previous step, resulting in a coherent but exclusive progression along a single trajectory. Although soft aggregation improves information capacity, it primarily reinforces the dominant path while introducing extraneous noise over time, as evidenced in Figures 1(c) and 1(d).
This phenomenon stems from the greedy nature of LLMs in autoregressive generation, where hidden states evolve at each step toward the most confident semantic direction [wu2025llms]. Such locally optimal decisions are amplified over long reasoning chains, resulting in a cumulative bias toward dominant paths and exclusion of alternatives. Thus, when a reasoning trajectory begins with a flawed premise, the lack of timely corrective feedback allows erroneous logic to propagate unimpeded through subsequent steps [xiong2025enhancing].
@@
This paradigm promotes robust policy optimization by allowing trajectories to benefit from cross-path insights. Specifically, the parameter-free gate facilitates step-level information exchange, providing trajectories with opportunities to gently refine their reasoning chain and reduce the persistence of local biases. In particular, by maintaining and refining multiple hypotheses in parallel, M3PO aligns with human-like cognition [fedorenko2024language, fedorenko2016language, benn2023language], striking an adaptive balance between exploration and exploitation.
@@ 2.1 Continuous Space Reasoning
Recent works have explored continuous space reasoning as an alternative to explicit CoT generation [pfau2024let, goyal2023think, wang2023guiding, wang2023towards, turpin2023language]. Several studies demonstrate that Transformers can perform multi-hop reasoning implicitly through their internal hidden states, without generating intermediate textual steps [yang2024large, shalev2024distributional]. Deng et al. [deng2023implicit, deng2024explicit] further exploit this idea by using hidden dynamics directly for reasoning, bypassing symbolic output altogether. Geiping et al. [geiping2025scaling] extend this paradigm with a depth-recurrent architecture that reuses Transformer layers during inference, increasing computational depth per token and enabling internal processing.
Another line focuses on explicit continuous reasoning through latent representations [deng2023implicit, tack2025llm]. COCONUT [hao2024training] enables reasoning in the hidden state space by replacing discrete tokens with continuous vectors, eliminating the need for human-readable chains. CODI [shen2025codi] learns to align recurrent state updates via self-distillation, forming structured reasoning trajectories in the latent space. Although conceptual advances, they often struggle to scale to larger models and maintain performance on complex reasoning tasks.
@@
More recently, training-free approaches such as Soft Thinking [zhang2025soft, gozeten2025continuous, wu2025llms] and Mixture-of-Inputs (MoI) [zhuang2025text] have emerged, leveraging output distributions to align hidden states with input embeddings, facilitating iterative reasoning in continuous space. Building on this, HRPO [yue2025hybrid] introduces a hybrid policy that fuses discrete token embeddings with soft embeddings through a gating mechanism, enabling richer internal state evolution during reasoning.
@@ 2.2 Reinforcement Learning
Reinforcement learning (RL) has emerged as a powerful framework for enhancing language models through feedback-driven decision making. In natural language generation, RL enables models to learn from implicit human preferences by maximizing cumulative rewards [ouyang2022training]. This is typically achieved with policy gradient methods such as REINFORCE [sutton1999policy], which update the policy using sampled responses and their rewards. Actor-critic approaches like A2C [mnih2016asynchronous] reduce variance with learned baselines, while PPO [schulman2017proximal] improves stability through a clipped objective that limits policy changes and prevents performance collapse.
An alternative line of work uses preference-based optimization, where models are fine-tuned on pairwise response comparisons [shakya2023reinforcement, chaudhari2025rlhf, zhang2025survey, wang2022deep]. Methods like DPO [rafailov2023direct] bypass explicit reward modeling by directly optimizing for preferred outputs, often eliminating the need for a reference model. Yet despite their efficiency, offline approaches generally lag behind online RL in complex reasoning settings. Recent online methods such as RLOO [ahmadian2024back], GRPO [shao2024deepseekmath], and REINFORCE++ [hu2025reinforce++] address variance and scalability by computing baselines from multiple responses in the same batch. Leveraging group-level comparisons, they achieve more stable and memory-efficient training, making them well suited for long-horizon reasoning tasks.
@@ 3.3 Policy Optimization
Note that M3PO uses raw log probabilities in its policy gradient (Equation (8)) rather than likelihood ratios as in PPO [schulman2017proximal] or GRPO [shao2024deepseekmath], eliminating the need for ratio clipping under our conservative update schedule. Furthermore, each trajectory is used only once for gradient updates since the hybrid representations are intrinsically tied to $\theta$, maintaining strict on-policy training. This makes M3PO both lightweight and compatible with other RL optimizations.
@@ 4.1 Evaluation on Knowledge Benchmarks
For knowledge-intensive reasoning, we use five widely adopted open-domain and multi-hop QA datasets: Natural Questions (NQ) [kwiatkowski2019natural], TriviaQA [joshi2017triviaqa], HotpotQA [yang2018hotpotqa], 2WikiMultiHopQA (2WikiMQA) [ho2020constructing], and Bamboogle [press2022measuring]. Following HRPO [yue2025hybrid], we employ the E5-base embedding model [wang2022text] to retrieve the top three Wikipedia documents as context for each prompt. We train on a joint corpus formed by merging the NQ and HotpotQA training sets, and we report exact match results on each dataset’s official evaluation split. Detailed implementation settings are provided in the Appendix.
For the 1.5B and 3B model sizes [qwen2.5techreport], we evaluate M3PO against representative post-training and retrieval baselines, including Supervised Fine-Tuning (SFT), the RL methods PPO [schulman2017proximal] and GRPO [shao2024deepseekmath], Retrieval-Augmented Generation (RAG) [lewis2020retrieval], and Hybrid Reasoning Policy Optimization (HRPO) [yue2025hybrid], an RL-based approach built on soft thinking [zhang2025soft]. We also report results with the larger Qwen2.5-7B-Instruct [qwen2.5techreport] backbone using standard QA and retrieval pipelines, namely direct inference (QA), Chain-of-Thought (CoT) [wei2022chain], interleaving retrieval with CoT (IRCoT) [trivedi2023interleaving], Search-o1 [li2025search], and RAG [lewis2020retrieval]. The best result in each block of Table 1 is highlighted in bold.
@@ 4.2 Evaluation on STEM Benchmarks
For reasoning-intensive STEM evaluation, we employ five benchmarks spanning mathematical and scientific domains: GSM8k [cobbe2021training], MATH [hendrycks2021measuring], MATH500 [lightman2023let], MMLU-STEM [hendrycks2020measuring], and ARC-Challenge [clark2018think]. For training, we use the GSM8k training split for GSM8k, the MATH training split for MATH and MATH500, and a merged corpus that combines the auxiliary MMLU and ARC-C training sets for MMLU-ST and ARC-C. Detailed settings are provided in Appendix.
For models at the 1.5B and 3B scales [qwen2.5techreport], we compare against several strong baselines: SFT, standard RL methods like PPO [schulman2017proximal] and GRPO [shao2024deepseekmath], and the RL-based latent reasoning method HRPO [yue2025hybrid]. To further contextualize the performance of M3PO, we also include comparisons with larger models ($\geq$ 7B parameters) using few-shot CoT reasoning, specifically DeepSeekMath-7B [shao2024deepseekmath], Gemma-2-9B [team2024gemma], Qwen2.5-7B [qwen2.5techreport], and MAmmoTH2-7B [yue2024mammoth2].
@@ 6 Implementation
M3PO is a lightweight multi-path collaborative learning framework designed to be compatible with any LLM architecture. It introduces no additional trainable parameters and preserves the original inference efficiency of the base model. To enable efficient training under this framework, we integrate optimized kernel implementations from Unsloth [^1] https://github.com/unslothai/unsloth and apply low-rank adaptation (LoRA) [hu2022lora] for parameter-efficient fine-tuning. All M3PO experiments adopt the hyperparameter settings detailed in Table 3, which serve as our default unless otherwise specified.
@@
The prompt construction follows a consistent template during both training and evaluation. Each input begins with a system message instructing the model to carry out step-by-step reasoning before generating its final answer, followed by the user query. The full prompt is then formatted according to the model’s native chat template. In line with established practice [yue2025hybrid], we use the minimal delimiter #### to separate the model’s reasoning chain from its final answer. This delimiter tokenizes as a single unit, adding no sequence length overhead while providing a clear signal for transitioning from latent collaborative reasoning to autoregressive answer generation. To ensure reasoning integrity, a penalty mechanism assigns zero reward to completions containing repeated delimiters, regardless of answer correctness. This prevents early termination of the reasoning process and encourages complete reasoning chains. Full prompt examples for different task types, including system messages and representative queries, are provided in Figures 9, 10, and 11, respectively.
@@
We use greedy decoding and the standard inference pipeline of the base LLM to ensure reproducibility. For knowledge tasks, exact match scores are reported on validation and test splits following [jin2025search]. For mathematical reasoning benchmarks including GSM8K, MATH, and MATH-500, along with multiple-choice datasets (MMLU-STEM and ARC-Challenge), we follow the post-processing and scoring procedures defined in [yue2024mammoth2].